Physical Chemistry

A vessel at 1000 K contains CO2 with a pressure of 0.5 atm. Some of the CO2 is convertedinto CO on the addition of graphite. If the total pressure at equilibrium is 0.8 atm, thevalue of K is

nikhil
12 Years agoGrade

4 Answers

Sunil Kumar FP
11 Years ago
CO2 +C ------2CO
initial presasure=.5atm
final pressure on addition of graphite
CO2=.5-X
CO=2X
total pressure at equilibrium=.5+x
but the total pressure at equilibrium= .8atm
,5+x=.8atm
x=.3atm
the value of reaction constant,K=[CO]^2/[CO2]=.6^2/.3=1.2
Elsa
10 Years ago
:)/(0.5-0.3)] =1.8atm 2]=[0.62/[CO2=0.5-XCO=2X Total pressure at equilibrium= 0.5+xbut the total pressure at equilibrium= 0.8atm0.5+x=0.8atm; x=0.3atm.The value of reaction constant, K=[CO]22CO Initial presasure=0.5atm Final pressure on addition of graphite: CO ------+C2CO
payal kendre
8 Years ago
CO2 pressure decreases let it be `x`then increase in pressure due to CO after reaction is 2xtotal pressure = (0.5-x) + 2x = 0.8 ATMX=0.3 ATMthus pressure of CO2 = 0.2and CO= 0.6K = P^2 CO ÷ P CO2 = 1.8
Kushagra Madhukar
5 Years ago
Dear student,
Please find the solution to your problem below.
 
CO2 + C →  2CO
initial presasure = 0.5 atm
final pressure on addition of graphite
CO2 = 0.5 – x
CO = 2x
total pressure at equilibrium = 0.5 + x
but the total pressure at equilibrium = 0.8 atm
0.5 + x = 0.8 atm
x = 0.3 atm
Hence, at equilibrium pressure of CO2 = 0.2 atm    and    CO = 0.6 atm
the value of reaction constant,K=[CO]2/[CO2] = 0.62/0.3 = 1.2
 
Thanks and regards,
Kushagra